CAVITY RING-DOWN LASER ABSORPTION SPECTROSCOPY OF IrC

\underline{TONGMEI MA}, J. W-H. LEUNG and A. S-C. CHEUNG, Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong.

The absorption spectrum of IrC at wavelength between 445-500 nm has been investigated using the technique of laser vaporization/reaction with free-jet expansion and cavity ring-down laser absorption (CRLA) spectroscopy. IrC molecule was produced by reacting laser-ablated iridium atoms with methane seeded in argon. This wavelength region covers the (0,0), (1,0) and (2,0) bands of the L ^{2}Φ_{7/2} - X ^{2}Δ_{5/2} transition. Analysis of the spectra gives refined band origins, vibrational and rotational constants for the L ^{2}Φ_{7/2} level.